# Jaike van Twiller1    Yossiri Adulyasak2    Erick Delage2    Djordje Grbic1    Rune Møller Jensen1

### v3

@@ Abstract
Reinforcement learning (RL) has shown promise in solving various combinatorial optimization problems. However, conventional RL faces challenges when dealing with real-world constraints, especially when action space feasibility is explicit and dependent on the corresponding state or trajectory. In this work, we focus on using RL in container shipping, often considered the cornerstone of global trade, by dealing with the critical challenge of master stowage planning. The main objective is to maximize cargo revenue and minimize operational costs while navigating demand uncertainty and various complex operational constraints, namely vessel capacity and stability, which must be dynamically updated along the vessel’s voyage. To address this problem, we implement a deep reinforcement learning framework with feasibility projection to solve the master stowage planning problem (MPP) under demand uncertainty. The experimental results show that our architecture efficiently finds adaptive, feasible solutions for this multi-stage stochastic optimization problem, outperforming traditional mixed-integer programming and RL with feasibility regularization. Our AI-driven decision-support policy enables adaptive and feasible planning under uncertainty, optimizing operational efficiency and capacity utilization while contributing to sustainable and resilient global supply chains.
1 Introduction
In recent years, machine learning (ML) for combinatorial optimization (CO) has gained much traction Bengio et al. (2021). Learning advanced planning policies yields promising results in solving CO problems in transportation and logistics. Such approaches can outperform existing solution methods in some well-known problems, such as vehicle routing Kool et al. (2019); Hottung et al. (2022) or job shop scheduling Kwon et al. (2020). Even though it is important to benchmark algorithms on traditional CO problems, a significant opportunity remains to address lesser-known yet critical planning challenges in supply chains. By addressing these challenges, we bridge the gap between theory and practice, enhancing the reliability and efficiency of worldwide supply chain operations.
Container shipping is a key component of the global supply chain, responsible for transporting 45% of annual goods valued at \$8.1 trillion UNCTAD (2021). Due to its scale, it is often regarded as the cornerstone of worldwide trade and modern consumerism. However, this scale also has a significant environmental impact, with yearly CO2 emissions exceeding 200 million tonnes Lloyd’s List (2022). Despite this, container vessels are considered an environmentally friendly mode of transportation, due to their relatively low emissions per cargo ton-mile Jensen et al. (2018). Consequently, container shipping is recognized to play a crucial role in the global green transition European Commission (2023). Within container shipping, there are several interdependent and complex planning tasks, such as berth allocation Martin-Iradi et al. (2022), pre-marshalling Hottung et al. (2020), quay crane scheduling Herup et al. (2022), and container vessel stowage planning Van Twiller et al. (2024b). Due to decision-making under demand uncertainty and various combinatorial aspects - such as capacity limits, seaworthiness requirements, minimizing overstowage and maximizing revenue - stowage planning is particularly challenging. To mitigate this challenge, stowage planning is often decomposed into the master planning problem (MPP), which involves the assignment of cargo to clusters of slots, and the slot planning problem (SPP), which allocates containers to individual slots Pacino et al. (2011). However, these subproblems in representative form, particularly the MPP, remain non-trivial and difficult to solve especially in the presence of uncertainty Van Twiller et al. (2024b).
In addition to its complexity, stowage planning faces public data scarcity and relies heavily on human planners supported by limited decision support systems Jensen et al. (2018). Planners should maximize capacity utilization while managing demand uncertainty by accounting for multiple future scenarios. However, evaluating such scenarios is computationally intractable, due to the problem’s complexity and dynamic nature. These limitations highlight the need for efficient, adaptive, and feasible decision-support systems, offering an opportunity for AI-driven solution methods to enhance the resilience and sustainability of the global supply chain.
This paper presents a deep reinforcement learning (DRL) approach with feasibility projection to construct adaptive and feasible solutions, providing a decision-support policy for the MPP under demand uncertainty.
Our main contributions are as follows:
  • Master Stowage Planning Environment: We develop a novel Markov decision process (MDP) for master stowage planning under demand uncertainty, incorporating realistic problem-specific constraints. To address data scarcity, we release the environment as an open-source implementation [^2] https://github.com/OptimalPursuit/navigating_uncertainty_in_mpp.
  • Feasibility Projection: We incorporate differentiable projection layers, including weighted scaling, policy clipping, and violation projection, to enforce inequality constraint satisfaction in DRL frameworks.
  • Efficient and Adaptive Solutions: Our experiments demonstrate that our policy efficiently generates adaptive and feasible solutions under demand uncertainty, significantly outperforming well-known DRL methods and a multi-stage stochastic MIP model.
  • Decision Support for Stowage Planning: Our decision-support policy transcends deterministic models, enabling dynamic and uncertainty-informed planning in a critical part of the global supply chain.
2 Domain Preliminaries
@@
We present the preliminaries of the stowage planning domain to provide the planning context. We refer to Jensen et al. (2018) for a comprehensive review.
@@
Voyages. Liner shipping companies utilize fleets of container vessels that operate on fixed schedules along a closed-loop course, similar to a maritime bus service that carries goods between ports. The complete journey covering all ports is called a voyage, typically starting in regions with a supply surplus (e.g., Asia) and navigating towards high-demand discharge areas (e.g., Europe), often voyages cover more than ten ports. A voyage can be described as a directed path graph $G_{P}=(P,E_{P})$ with nodes being $P=\{1,2,\dots,N_{P}\}$, and edges being legs between ports $E_{P}=\{(p,p\!+\!1)\mid p\in\{1,2,\dots,N_{P}\!-\!1\}\}$ with $N_{P}$ being the last port. We also define sub-voyages by set ${P}_{\text{start}}^{\text{end}}=\{p\in{P}\!\mid\!\text{start}\leq p\leq\text{end}\}$.
Containers have a port of load (pol) and a port of discharge (pod), also named transports or origin-destination pairs. Consider $\textit{pol}\in{P}_{1}^{N_{P}-1}$ and $\textit{pod}\in{P}_{2}^{N_{P}}$ with $P^{2}=P\times P$, which can be represented by transport $\textit{tr}=(\textit{pol},\textit{pod})$. We can define a set of transports $\textit{TR}=\{(i,j)\in{P}^{2}\!\mid\!i<j\}$.
[…]
@@
Vessel Layout. Figure 1 shows the cellular layout of container vessels with standardized coordinates of (bay, row, tier). It is divided into bays (02-38) ranging from the front (e.g., bay 02) to the rear (e.g., bay 38), also named from fore to aft. Bays are defined by an ordered set ${B}=\{1,2,\dots,N_{B}\}$, ranging from fore to aft with $N_{B}$ being the last bay. Each bay contains stacks of cells, whereas cells contain two slots that can hold two 20-foot containers or one 40-foot container. Bays are horizontally separated by hatch covers into above and below deck sections, introducing the set of decks ${D}=\{d^{\textit{above}},d^{\textit{below}}\}$. It is worth noting that hatch covers introduce complexity since below-deck cargo can only be discharged if the above-deck cargo is cleared.
@@
Utilization. Vessel utilization $u_{p}\in\mathbb{Z}^{|B|\times|D|\times|K|\times|\textit{TR}|}$ represents the cargo placement across bays $B$, decks $D$, cargo types $K$, and transports TR. Load operations are defined by $u^{+}_{p}\in\mathbb{Z}_{>0}^{|B|\times|D|\times|K|\times|\textit{TR}|}$, whereas discharging is denoted by $u^{-}_{p}\in\mathbb{Z}_{>0}^{|B|\times|D|\times|K|\times|\textit{TR}|}$. The utilization at port $p$ is defined as $u_{p}=u_{p-1}+u^{+}_{p}-u^{-}_{p}\;\forall p\in{P}$, where $u_{0}$ is the vessel’s arrival condition at the first port. We also define the vessel’s pre-loading utilization $u^{\prime}_{p}=u_{p-1}-u^{-}_{p}$.
Objectives and Global Constraints. The MPP aims to create a capacity management plan for each general location (indexed by bay and deck) while adhering to global constraints, as outlined below. The MPP maximizes revenue by optimizing capacity for high-value cargo and minimizes costs through improved operational efficiency. The SPP then assigns containers to slots within the general locations while considering local constraints. Due to frequent re-planning as new information arises, algorithmic runtimes over 10 minutes are impractical for real-world operations Pacino et al. (2011).
Vessels have different types of capacity related to TEUs, height, weight and special cargo. The amount to which this can be utilized depends on the arrival condition of vessels and uncertain cargo demand. It is therefore non-trivial to ensure optimal utilization of vessel capacities. The key efficiency objectives are minimizing overstowage, an NP-hard task Tierney et al. (2014), and minimizing the quay crane makespan. Overstowage occurs when containers stacked on top block access to those below in need of discharging, requiring additional crane operations to remove and reload the obstructing containers, thereby increasing the workload. In addition, the crane makespan considerably impacts the port stay duration, while ports set makespan targets and overruns incur costs. It is thus beneficial to minimize the makespan by maximizing the average use of parallel cranes, thereby minimizing idle time.
Furthermore and most importantly, vessels must adhere to safety requirements to ensure seaworthiness, which is heavily influenced by weight distribution. Before introducing the constraints, we first clarify that $\otimes$ represents the outer product and $\odot$ denotes the element-wise Hadamard product, with broadcasting applied when shapes are compatible. The longitudinal (lcg) and vertical centres of gravity (vcg), which are associated with trim and metacentric height, of the load $u_{p}$ prior to leaving port $p$ must remain within specified bounds ($\underline{\textit{lcg}},\overline{\textit{lcg}}\in\mathbb{R}_{>0}$), as shown in Constraints (1) and (2). To compute the stability, we define the longitudinal moment $\textit{lm}=\textit{ld}\otimes w$ and vertical moment $\textit{vm}=\textit{vd}\otimes w$, where $w\in\mathbb{R}^{|K|\times|\textit{TR}|}_{>0}$ represents container weights, and $\textit{ld},\textit{vd}\in\mathbb{R}^{|B|\times|D|}$ denote the longitudinal and vertical distances of positions from the vessel’s centre of gravity.
@@
In conclusion, the scale and complexity of the problem pose significant challenges. As a result, most models focus on subsets of realistic constraints to remain tractable, often requiring human intervention to ensure practical usability.
3 Related Work
This section positions our work in the literature on stowage planning, stochastic programming and ML in optimization.
@@ 3 Related Work
Stowage Planning. To address the complexity of stowage planning, various solution approaches have been applied to different problem formulations, including exact methods Roberti and Pacino (2018), linearly relaxed MIP Pacino et al. (2011), matheuristics Parreño-Torres et al. (2021), population-based metaheuristics Chang et al. (2022), neighbourhood-based metaheuristics Pacino (2018), hybrid frameworks Bilican et al. (2020). Despite these techniques, a recent survey highlights that scalable solutions for representative stowage and MPP problems remain unsolved Van Twiller et al. (2024b). Research has largely focused on deterministic problems, overlooking real-world needs for profit maximization in the face of demand uncertainty.
@@
Stochastic Programming. Decision-making under uncertainty is traditionally approached through stochastic programming, where uncertainty is explicitly represented as a set of discrete scenarios evolving over multiple stages in time, approximating the underlying probability distribution Birge and Louveaux (2011). This process results in a scenario tree, whose computational complexity grows exponentially as $\mathcal{O}(b^{T})$, where $b$ is the branching factor and $T$ is the number of stages. Hence, multi-stage problems with $T>2$ and a reasonably large $b$ are often intractable. Common techniques to overcome this complication are scenario reduction to reduce problem size Römisch (2009), decomposition techniques like Benders decomposition or Lagrangian relaxation to divide the problem Rahmaniani et al. (2017), progressive hedging to enforce non-anticipativity iteratively Boland et al. (2018), or approximation methods, such as stochastic dual dynamic programming Shapiro (2011) and sample average approximation Chen and Luedtke (2022), to enhance tractability while preserving solution quality.
Learning with Hard Constraints. In learning solution heuristics, a challenge arises to ensure feasibility in complex action spaces. This challenge is exacerbated when dealing with the dependence on state variables to define the feasible region. Several deep learning approaches have dealt with learning constraints by backpropagation through (in)equality completion Donti et al. (2021), or differentiable projection layers, such as mapping interior points to boundaries Li et al. (2023), convex programming layers Agrawal et al. (2019), or problem-specific repair layers Chen et al. (2024). Moreover, safe reinforcement learning has dealt with constraints by, e.g., constrained MDPs with a primal-dual approach Ding et al. (2020), soft barriers function Wang et al. (2023), and safety shields Alshiekh et al. (2018).
4 Markov Decision Processes
We present a formal MDP and a decomposed MDP to solve using DRL approaches. Technical details about both MDPs are provided in Appendix A.
4.1 Formal MDP
@@
Let ${\mathcal{M}}=({S},{X},\mathcal{T},{\mathcal{R}},{P}^{N_{P}-1}_{1},\gamma)$ define an episodic discounted MDP representing the MPP, where ${S}$ is a set of states, ${X}$ is a set of actions, $\mathcal{T}:{S}\times{X}\rightarrow\Delta({S})$ is the transition function, ${\mathcal{R}}:{S}\times{X}\times{P}^{N_{P}-1}_{1}\rightarrow\mathbb{R}$ is the reward function, ${P}^{N_{P}-1}_{1}$ is the finite horizon of load ports with $N_{P}-1$ being the last port, and $\gamma\in(0,1]$ is the discounting factor.
State. The state is given by ${s}_{p}\in{S}$, defined as ${s}_{p}=(u_{p},q_{p},\zeta)$. This includes vessel utilization $u_{p}\in\mathbb{R}_{\geq 0}^{n_{u}}$ and realized demand $q_{p}\in\mathbb{R}_{\geq 0}^{n_{q}}$, where $n_{u}=|B|\times|D|\times|K|\times|\textit{TR}|$, $n_{c}=|B|\times|D|$ and $n_{q}=|K|\times|\textit{TR}|$ are the shapes of utilization, location and demand, respectively.
Additionally, environment parameters $\zeta$ contain the expected value $\mu\in\mathbb{R}^{n_{q}}_{\geq 0}$ and standard deviation $\sigma\in\mathbb{R}^{n_{q}}_{>0}$ of demand, load ports $i$, discharge ports $j$ and cargo types $k$ as $(i,j,k)\in\textit{TR}\times K$, TEU per container $\textit{teu}\in\{1,2\}^{n_{q}}$, cargo weight $w\in\mathbb{R}^{n_{q}}_{>0}$, cargo revenue $\textit{rev}\in\mathbb{R}^{n_{q}}_{>0}$.
The initial state $s_{0}=(u_{0},q_{0},\zeta)$ consists of an empty vessel $u_{0}=\mathbf{0}^{n_{u}}$, realized demand $q_{0}$ at initial port, and $\zeta$ is initialized randomly for each episode.
Action. An action ${x}_{p}\in{X}$ assigns a real number of containers to utilization $u_{p}$, where ${x}_{p}\in\mathbb{R}_{>0}^{n_{u}}$ and thus similar to $u^{+}_{p}$. Each action ${x}_{p}$ is subject to a feasible region, defined by polyhedron $\textit{PH}({s}_{p})=\{{x}_{p}\in\mathbb{R}^{n_{u}}_{>0}:A({s}_{p}){x}_{p}\leq b({s}_{p})\}$. Here, $A({s}_{p})\in\mathbb{R}^{m_{u}\times{n_{u}}}$ is the constraint matrix, $b({s}_{p})\in\mathbb{R}^{m_{u}}$ is the bound vector, and $m_{u}$ is the number of constraints.
Previous work demonstrated that linearly relaxed MPPs outperform exact MPPs, as the subsequent SPP effectively discretizes the solution Pacino et al. (2011). Accordingly, we can adopt real-valued actions instead of integer-valued actions. Given a traditional MPP formulation with utilization $u_{p}$ Van Twiller et al. (2024a), we define constraints of $\textit{PH}({s}_{p})$ in terms of actions ${x}_{p}$ and pre-loading utilization $u^{\prime}_{p}$.
@@
-$\text{-}\mathbf{1}^{\top}\!\big{(}(\textit{lm}\!-\!\underline{\textit{lcg}}w)\!\odot\!{x}_{p}\big{)}$$\leq\!\text{-}\mathbf{1}^{\top}\!\big{(}(\underline{\textit{lcg}}w\!-\!\textit{lm})\!\odot\!u^{\prime}_{p}\big{)}$ (5)
@@
-$\text{-}\mathbf{1}^{\top}\!\big{(}(\textit{vm}\!-\!\underline{\textit{vcg}}w)\!\odot\!{x}_{p}\big{)}$$\leq\!\text{-}\mathbf{1}^{\top}\!\big{(}(\underline{\textit{vcg}}w\!-\!\textit{vm})\!\odot\!u^{\prime}_{p}\big{)}$ (7)
@@
Constraint (3) limits load ${x}_{p}$ to the available demand $q_{p}$, while Constraint (4) ensures ${x}_{p}$ does not exceed the residual TEU capacity. Constraints (5)–(6) enforce lcg limits, while Constraints(7)–(8) impose similar bounds on the vcg. These stability bounds are derived from Constraints (1) and (2).
Transition. We use a stochastic transition function $\mathcal{T}({s}_{p+1}|{s}_{p},{x}_{p})\in\Delta({S})$. Within an episode, the transition consists of multiple components:
  • Upon port arrival, port demand $q_{p}$ is revealed.
  • Subsequently, onboard cargo is discharged $u_{p+1}=u_{p}\odot(1-\mathbf{e}_{p}^{-})$, where $\mathbf{e}_{p}^{-}\in\{0,1\}^{n_{q}}$ is a binary mask indicating the cargo type and transport to nullify in $u_{p}$.
  • Finally, cargo is loaded onboard $u_{p+1}=u_{p}+{x}_{p}$. Action ${x}_{p}$ is based on the current utilization $u_{p}$ and revealed demand $q_{p}$ of port $p$. Future port demand stays unknown.
Reward. Equation (4.1) defines our deterministic reward function, computing profit as the difference between revenue and costs. Revenue is computed as the sum of $x_{p}$ corresponding to elements of $q_{p}$. Since revenue cannot be obtained from containers exceeding demand, the summation of $x_{p}$ is restricted to the elements within $q_{p}$. While costs are computed via the state-dependent auxiliary variables hatch overstows $\textit{ho}({s}_{p},p)\in\mathbb{R}^{|B|}_{>0}$ and excess crane moves $\textit{cm}({s}_{p},p)\in\mathbb{R}^{|B|-1}_{>0}$. These costs are weighted by $\textit{ct}^{\textit{ho}}\in\mathbb{R}_{>0}$ and $\textit{ct}^{\textit{cm}}\in\mathbb{R}_{>0}$, respectively.

### v5

# Navigating Demand Uncertainty in Container Shipping: Deep Reinforcement Learning for Enabling Adaptive and Feasible Master Stowage Planning

@@ Abstract
Reinforcement learning (RL) has shown promise in solving various combinatorial optimization problems. However, conventional RL faces challenges when dealing with complex, real-world constraints, especially when action space feasibility is explicit and dependent on the corresponding state or trajectory. In this work, we address stochastic sequential dynamic decision-making problems with state-dependent constraints. As a relevant and real-world case study, we focus on the master stowage planning problem in container shipping, which aims to optimize revenue and operational costs under demand uncertainty and operational constraints. We propose a deep RL framework with an encoder-decoder model and feasibility layers that satisfy convex constraints and maintain unbiased gradient flow, which embed problem instances, current solutions, and demand uncertainty to guide learning. Experiments show that our model efficiently finds adaptive, feasible solutions that generalize across varying distributions and scale to larger instances, outperforming state-of-the-art baselines in constrained RL and stochastic programming. By uniting artificial intelligence and operations research, our policy empowers humans to make adaptive, uncertainty-aware decisions for resilient and sustainable planning.
Introduction
In recent years, machine learning (ML) for combinatorial optimization (CO) has gained significant traction, driving advances in solving complex decision problems (Bengio, Lodi, and Prouvost 2021). Learning advanced planning policies yields promising results in solving established problems, such as vehicle routing (Kool, van Hoof, and Welling 2019; Hottung, Kwon, and Tierney 2022) or job shop scheduling (Kwon et al. 2020). Although benchmarking algorithms on well-known problems is essential, much work on ML for CO focuses on simplified, deterministic, and static settings with implicit feasibility. In contrast, real-world problems are often stochastic, dynamic, and subject to explicit constraints dependent on states or trajectories. We address this challenge by bridging artificial intelligence (AI) and operations research for practical and impactful planning.
A relevant domain is container shipping, which transports 45% of global goods worth \$8.1 trillion annually (United Nations Conference on Trade and Development 2021), emits over 200 million tonnes of CO2 yearly (Lloyd’s List 2022), yet remains key to the green transition due to its relatively low emissions per ton-mile (European Commission 2023). Container shipping has highly combinatorial, uncertain, and operationally constrained problems, i.e., berth allocation (Martin-Iradi, Pacino, and Ropke 2022), pre-marshalling (Hottung, Tanaka, and Tierney 2020), quay crane scheduling (Herup et al. 2022), and container vessel stowage planning (Van Twiller et al. 2024b). The latter can decompose into two persistently challenging subproblems: the master planning problem (MPP) assigns cargo to clusters of slots, and the slot planning problem (SPP) allocates containers to individual slots (Pacino et al. 2011). These problems involve stochastic, dynamic decisions subject to constraints, where actions affect future feasibility and poor decisions cause disruptions and unnecessary emissions in global supply chains.
This paper presents a deep reinforcement learning (DRL) framework with feasibility projection that learns to act in complex, uncertain environments while enforcing hard constraints. Conventional RL struggles with feasibility, while stochastic programs and implicit optimization incur high computational costs. Instead, our method efficiently learns from experience and corrects infeasible actions. We formulate the MPP under demand uncertainty as a Markov decision process (MDP) and solve it using a projected encoder-decoder model that integrates instance, solution, and uncertainty data. Differentiable projections with Jacobian corrections ensure unbiased gradient flow and enable stable training of feasible policies. Our main contributions are:
  • MPP Environment: A novel stochastic, dynamic MDP with state-dependent constraints for the MPP under demand uncertainty to benchmark AI policies and address data scarcity in shipping. Released as open-source code [^1] https://tinyurl.com/navigating-mpp-uncertainty
  • Unbiased Violation Projection: We propose differentiable feasibility projections that ensure satisfaction of general convex and problem-specific constraints, including unbiased gradient flow through Jacobian corrections.
  • Scalable and Adaptive Solutions: Our model uniquely surpasses state-of-the-art constrained RL and stochastic MIP baselines by efficiently generating feasible, high-profit solutions that scale and adapt to varied instances.
  • Practical Decision Support: We demonstrate how AI policies assist human planners in making uncertainty-aware decisions to address real-world challenges.
Domain Preliminaries
@@ Introduction
We present some domain preliminaries to provide context. For a comprehensive review, we refer to Jensen et al. (2018).
Voyages and Container Transport.
@@
Liner shipping companies operate vessels on fixed-schedule, closed-loop voyages, much like a maritime bus service. These journeys, often exceeding ten ports, typically originate in supply-surplus regions (e.g., Asia) and finish in high-demand areas (e.g., Europe). A voyage can be described as a directed path graph $G_{P}=(P,E_{P})$, where $P=\{1,2,\dots,N_{P}\}$ are the nodes representing ports, and $E_{P}=\{(p,p\!+\!1)\mid p\in\{1,2,\dots,N_{P}\!-\!1\}\}$ are the edges representing legs between ports. We define sub-voyages by the set ${P}_{\text{start}}^{\text{end}}=\{p\in{P}\!\mid\!\text{start}\leq p\leq\text{end}\}$. Containers have a port of load (pol) and a port of discharge (pod), also named transports or origin-destination pairs. A transport $\textit{tr}=(\textit{pol},\textit{pod})$ is defined with $\textit{pol}\in{P}_{1}^{N_{P}-1}$ and $\textit{pod}\in{P}_{2}^{N_{P}}$. The set of all possible transports is $\textit{TR}=\{(i,j)\in{P}^{2}\!\mid\!i<j\}$.
Container Cargo.
Container cargo is categorized by key characteristics: lengths (20 ft, 40 ft), weight classes (light, medium, heavy), and customer contracts (long-term, spot market) that determine cargo revenue (rev). Long-term contracts offer lower but stable revenue and high volumes, taking priority over spot market contracts, which provide higher yet volatile revenue in lower volumes. The lack of no-show fees creates uncertainty in container demand, although this becomes more predictable closer to the arrival date. This necessitates dynamic cargo allocation to vessels during the voyage. Containers are often grouped into cargo classes by characteristics, such as ${K}=\{20\textit{ft},40\textit{ft}\}\times\{\textit{Light},\textit{Medium},\textit{Heavy}\}\times\{\textit{Spot},\textit{Long}\}$.
Vessel Layout.
@@
Figure 1 shows the cellular layout of container vessels with standardized coordinates of (bay, row, tier). It is divided into bays (02-38) ranging from the front (e.g., bay 02) to the rear (e.g., bay 38), also named from fore to aft. Bays are defined by an ordered set ${B}=\{1,2,\dots,N_{B}\}$, ranging from fore to aft with $N_{B}$ being the last bay. Each bay contains stacks of cells, whereas cells contain two slots that can hold two 20-foot containers or one 40-foot container. Bays are horizontally separated by hatch covers into above and below deck sections, introducing the set of decks ${D}=\{d^{\textit{above}},d^{\textit{below}}\}$. It is worth noting that hatch covers introduce complexity since below-deck cargo can only be discharged if the above-deck cargo is cleared.
@@
Utilization.
Vessel utilization $u_{p}\in\mathbb{Z}^{|B|\times|D|\times|K|\times|\textit{TR}|}_{\geq 0}$ represents the cargo placement across bays $B$, decks $D$, cargo types $K$, and transports TR. Load operations are defined by $u^{+}_{p}\in\mathbb{Z}_{>0}^{|B|\times|D|\times|K|\times|\textit{TR}|}$, whereas discharging is denoted by $u^{-}_{p}\in\mathbb{Z}_{>0}^{|B|\times|D|\times|K|\times|\textit{TR}|}$. The utilization at port $p$ is defined as $u_{p}=u_{p-1}+u^{+}_{p}-u^{-}_{p}\;\forall p\in{P}$, where $u_{0}$ is the vessel’s arrival condition at the first port. We also define the vessel’s pre-loading utilization $u^{\prime}_{p}=u_{p-1}-u^{-}_{p}$.
Objectives and Global Constraints.
The MPP creates a capacity management plan for each general location (bay, deck) to maximize revenue and minimize costs, while the subsequent SPP assigns containers to specific slots. Due to multi-port voyages, each decision affects future feasibility and optimality. Due to frequent re-planning, runtimes over 10 minutes are impractical (Pacino et al. 2011).
Key efficiency objectives include minimizing overstowage (an NP-hard problem (Tierney, Pacino, and Jensen 2014)) and crane makespan. Overstowage occurs when containers block access to those below needing discharge, requiring unnecessary moves. Minimizing makespan reduces port stay duration and associated costs.
Furthermore, vessels must satisfy safety constraints for seaworthiness, particularly their weight distribution. The longitudinal (lcg) and vertical (vcg) centers of gravity for load $u_{p}$ must remain within bounds, where $\otimes$ denotes outer product and $\odot$ element-wise product:
@@
where $\textit{lm}=\textit{ld}\otimes w$ and $\textit{vm}=\textit{vd}\otimes w$ are longitudinal and vertical moments, with $w\in\mathbb{R}^{|K|\times|\textit{TR}|}_{>0}$ being container weights, and $\textit{ld},\textit{vd}\in\mathbb{R}^{|B|\times|D|}$ denoting the respective longitudinal and vertical distances from the center of gravity.
The problem’s scale and complexity, combined with data scarcity and runtime limits, make most models focus on constraint subsets for tractability. This highlights the need for efficient AI-driven solutions to enhance decision support.
Related Work
@@
Various solution methods have been proposed to address container stowage planning, including exact algorithms (Roberti and Pacino 2018), relaxed MIPs (Pacino et al. 2011), matheuristics (Parreño-Torres et al. 2021), metaheuristics (Chang, Hamedi, and Haghani 2022; Pacino 2018), and hybrid frameworks (Bilican, Evren, and Karatas 2020). A recent survey highlights that scalable solutions for representative stowage planning and the MPP remain unsolved (Van Twiller et al. 2024b). Crucially, most research assumes deterministic problems with cost minimization, overlooking real-world needs for profit maximization in the face of demand uncertainty. Our work targets this gap by optimizing the MPP under demand uncertainty.
Stochastic Programming.
@@
Traditionally, uncertainty in optimization is handled using stochastic programming (Birge and Louveaux 2011). It models discrete scenarios over multiple stages, forming a scenario tree with exponential complexity $\mathcal{O}(b^{T})$, where $b$ is the branching factor and $T$ the number of stages. Consequently, multi-stage problems with $T>2$ and a reasonably large $b$ are often intractable. Common mitigations include scenario reduction (Römisch 2009), Benders or Lagrangian decomposition (Rahmaniani et al. 2017), and progressive hedging to enforce non-anticipativity iteratively (Boland et al. 2018). Alternatively, approximation methods improve tractability: stochastic dual dynamic programming applies temporal decomposition but depends on known uncertainty distributions (Shapiro 2011), while sample average approximation uses Monte Carlo sampling to solve a deterministic surrogate (Chen and Luedtke 2022). Model-free RL bypasses most of the mentioned drawbacks by learning policies through direct interaction, implicitly modeling uncertainty for dynamic environments (Powell 2022). However, it faces challenges in sample efficiency and handling complex constraints.
Learning with Hard Constraints.
Enforcing constraints in differentiable models is challenging because it requires balancing feasibility and objectives. Soft feasibility regularization, such as penalty functions (Zhang et al. 2022), barrier methods (Wang et al. 2023), or primal-dual algorithms (Ding et al. 2020), lacks strict guarantees and needs careful tuning. Decoupled methods, e.g., post-hoc repair or safety shields (Alshiekh et al. 2018), ensure feasibility but sacrifice differentiability, limiting end-to-end learning. Integrated techniques, like differentiable implicit optimization (Agrawal et al. 2019) or differentiable projection layers (Donti, Rolnick, and Kolter 2021; Chen, Tanneau, and Van Hentenryck 2024), embed constraints directly but produce biased gradients and computational overhead. Our work introduces a differentiable violation projection that enforces hard convex constraints and ensures unbiased gradients via Jacobian correction, enabling efficient learning in constrained problems.
Markov Decision Processes
We formulate the MPP as an MDP with action constraints. Intuitively, the MDP tracks what’s loaded and needed at each port (state), chooses how to load (action), updates the vessel and demand per port (transition), and measures profit after every decision (reward). To improve learning efficiency, we decompose the MDP into a sequence of simpler decisions, thereby effectively reducing the action space.
Markov Decision Process
@@
Let ${\mathcal{M}}=({S},{X},\mathcal{T},{\mathcal{R}},{P}^{N_{P}-1}_{1},\gamma)$ define an episodic discounted MDP representing the MPP, where S{S} is a set of states, X{X} is a set of actions, $\mathcal{T}:{S}\times{X}\rightarrow\Delta({S})$ is the transition function, ${\mathcal{R}}:{S}\times{X}\times{P}^{N_{P}-1}_{1}\rightarrow\mathbb{R}$ is the reward function, ${P}^{N_{P}-1}_{1}$ is the finite horizon of load ports with last port $N_{P}-1$, and $\gamma\in(0,1)$ is the discounting factor.
State.
The state ${s}_{p}=(u_{p},q_{p},\zeta)$ includes vessel utilization $u_{p}\in\mathbb{R}_{\geq 0}^{n_{u}}$, realized demand $q_{p}\in\mathbb{R}_{\geq 0}^{n_{q}}$, and instance parameter $\zeta$ (such as demand statistics, port and cargo indices, TEU, weight, revenue; details in Appendix A). The initial state $s_{0}=(u_{0},q_{0},\zeta)$ has an empty vessel $u_{0}=\mathbf{0}^{n_{u}}$, realized demand $q_{0}$ of first port, and a generated problem instance $\zeta$. Note that $n_{u}=|B|\times|D|\times|K|\times|\textit{TR}|$, $n_{c}=|B|\times|D|$ and $n_{q}=|K|\times|\textit{TR}|$ are the shapes of utilization, location and demand, respectively.
Action.
At each port $p$, an action $x_{p}\in\mathbb{R}_{\geq 0}^{n_{u}}$ specifies how many containers of each group to load at each vessel position. To ensure safe and efficient loading, actions must satisfy a set of linear constraints defined by the polyhedron $\textit{PH}(s_{p})=\left\{x_{p}\in\mathbb{R}^{n_{u}}_{\geq 0}:A(s_{p})x_{p}\leq b(s_{p})\right\},$ where $A(s_{p})\in\mathbb{R}^{m_{u}\times n_{u}}$ is the constraint matrix and $b(s_{p})\in\mathbb{R}^{m_{u}}$ is the vector of bounds, with $m_{u}$ different constraints.
The constraints below are expressed in terms of the action $x_{p}$ and the pre-loading utilization $u^{\prime}_{p}$, meaning the vessel’s state after discharging cargo at the port. We use real-valued actions, as downstream planning stages (e.g., the SPP) will discretize these plans (Pacino et al. 2011).
Constraint (3) limits load ${x}_{p}$ to the available demand $q_{p}$, while Constraint (4) ensures ${x}_{p}$ does not exceed the residual TEU capacity. Constraints (5)–(6) enforce lcg limits, while Constraints (7)–(8) impose similar bounds on the vcg. These stability bounds are derived from Constraints (1) and (2). Full derivations of $\textit{PH}(s_{p})$ are provided in Appendix A.
@@
-$\text{-}\mathbf{1}^{\top}\!\big{(}(\textit{lm}\!-\!\underline{\textit{lcg}}w)\!\odot\!{x}_{p}\big{)}$ -$\leq\!\text{-}\mathbf{1}^{\top}\!\big{(}(\underline{\textit{lcg}}w\!-\!\textit{lm})\!\odot\!u^{\prime}_{p}\big{)}$ (5)
@@
-$\text{-}\mathbf{1}^{\top}\!\big{(}(\textit{vm}\!-\!\underline{\textit{vcg}}w)\!\odot\!{x}_{p}\big{)}$ -$\leq\!\text{-}\mathbf{1}^{\top}\!\big{(}(\underline{\textit{vcg}}w\!-\!\textit{vm})\!\odot\!u^{\prime}_{p}\big{)}$ (7)
@@
Transition.
We use a stochastic transition function $\mathcal{T}(s_{p+1}\mid s_{p},x_{p})$ to update the state at each port. Each step of the transition, when moving to port $p+1$, includes:
  • Upon port arrival, the true demand $q_{p+1}$ is revealed, while future port demand remains unknown.
  • Cargo destined for this port is unloaded: $u_{p+1}=u_{p}\odot(1-\mathbf{e}_{p}^{-})$, with $\mathbf{e}_{p}^{-}$ marking which cargo to discharge.
  • New cargo is loaded: $u_{p+1}=u_{p+1}+x_{p}$, where action $x_{p}$ should fit feasible region $\textit{PH}(s_{t})$.
Reward.
Equation (Reward.) defines a deterministic reward function that computes profit as revenue minus costs. Revenue is calculated as the sum of loaded containers $x_{p}$ that match actual demand $q_{p}$; no revenue is received for containers exceeding demand. Costs are determined using two state-dependent auxiliary variables: hatch overstows $\textit{ho}(s_{p},p)\in\mathbb{R}^{|B|}_{>0}$ and excess crane moves $\textit{cm}(s_{p},p)\in\mathbb{R}^{|B|-1}_{>0}$. Both components are weighted, respectively, by coefficients $\textit{ct}^{\textit{ho}}\in\mathbb{R}_{>0}$ and $\textit{ct}^{\textit{cm}}\in\mathbb{R}_{>0}$.